### v5

@@ Abstract
In recent decades, several assistive technologies have been developed to improve the ability of blind and visually impaired individuals to navigate independently and safely. At the same time, simultaneous localization and mapping (SLAM) techniques have become sufficiently robust and efficient to be adopted in developing these assistive technologies. We present the first systematic literature review of 54 recent studies on SLAM-based solutions for blind and visually impaired people, focusing on literature published from 2017 onward. This review explores various localization and mapping techniques employed in this context. We systematically identified and categorized diverse SLAM approaches and analyzed their localization and mapping techniques, sensor types, computing resources, and machine-learning methods. We discuss the advantages and limitations of these techniques for blind and visually impaired navigation. Moreover, we examine the major challenges described across studies, including practical challenges and considerations that affect usability and adoption. Our analysis also evaluates the effectiveness of these SLAM-based solutions in real-world scenarios and user satisfaction, providing insights into their practical impact on BVI mobility. The insights derived from this review identify critical gaps and opportunities for future research activities, particularly in addressing the challenges presented by dynamic and complex environments. We explain how SLAM technology offers the potential to improve the ability of visually impaired individuals to navigate effectively. Finally, we present future opportunities and challenges in this domain.
@@ II-B3 Selection criteria
Table V presents the selection criteria used to identify the eligible studies during the selection process. The Availability criterion included studies accessible in full text from digital databases. In addition, the Language criterion ensured the inclusion of publications written only in English. Furthermore, the Publication Period criterion restricted the inclusion of studies published between January 2017 and July 2023. This timeframe allowed us to prioritize the most current and state-of-the-art approaches. The Type of Source criterion included conference and journal papers, which were considered peer-reviewed and academically recognized sources. Books, dissertations, newsletters, speeches, technical reports, and white papers were excluded. Finally, the Relevance criterion played an important role in the exclusion process; therefore, publications outside the scope of our study were excluded based on a review of their titles and abstracts.
@@ III Result
To further understand the technical features employed in these solutions, the detailed information is presented in Tables XVIIXIX. These tables focus on the localization and mapping components of the assistive system, specifically highlighting the sensor types, computing resources, and application of machine learning-based methods. By examining these features, we can gain deeper insight into how these systems are structured and the diverse technologies utilized to achieve accurate and efficient SLAM for assistive navigation. It is important to note that this information relates only to the localization and mapping components of the assistive navigation solutions. Details of the entire system are provided in Tables XXXIIIXXXVII.
@@ III-A1 Approaches
Dynamic-SLAM Feature-based Optimization-based Visual [90]
@@ III-A3 Computing resource
To process data and run localization and mapping algorithms, the reviewed studies adopted two classes of computational resources: local and remote. Local computations are performed in situ on devices, such as smartphones, tablets, laptops, portable microcontrollers, and UP board computers. In some cases, algorithms were applied on PCs. Table XX categorizes the computing resources used in the reviewed studies for localization and mapping tasks. Information regarding the computing resources for the entire navigation assistive system is detailed in Tables XXXV-XXXVII
TABLE XX: Categorization of computing resources used for localization and mapping tasks.
@@ III-B1 Advantages
The cost-effectiveness of SLAM-based solutions attributed to their reliance on widely available low-cost sensors makes this technology particularly interesting. Systems employing monocular cameras or wearable RGB-D cameras exemplify how SLAM can be implemented in a cost-effective manner without compromising functionality, thus making advanced navigational aids accessible to more users [69, 85]. The advantages of SLAM, derived from the literature, are listed in Table XXI.
TABLE XXI: Key features of SLAM that benefit visually impaired navigation.
@@ III-B2 Limitations
Some SLAM applications require external calibration or setup, such as placement of calibration boards in specific environments. This requirement can limit the spontaneity and ease of use of SLAM-based navigation aids because it imposes additional constraints [105]. Table XXII outlines the overall limitations of SLAM, derived from the publications under review.
TABLE XXII: Limitations of SLAM for visually impaired navigation.
@@ III-C1 Technical and methodological challenges
Optimal pathfinding, perception of surroundings, and obstacle avoidance are crucial for navigation. Precise localization of a visually impaired user within the environment is essential for the effective operation of these functions. Our surveyed papers addressed the localization and mapping problems using various techniques. Some of these studies explored other navigation-related challenges. We categorized these challenges into two groups: those relevant to environmental complexities, and those related to the sensors used to receive environmental data. Dynamic obstacles and crowded spaces constitute the challenges in the first group, whereas challenges related to changing lighting conditions and the rapid motion of users that results in motion blur fall into the second group. Table XXIII lists studies that investigated these challenges through the integration of SLAM with other approaches. In the following section, we discuss the studies that address these challenges.
TABLE XXIII: Challenges addressed in reviewed studies using SLAM techniques
@@
[59] presented a guide mobile robot engineered for the complexities of navigating different environments while considering dynamic objects and human presence. The robot could handle crowded environments with multiple dynamic objects. To accomplish this, the robot leveraged a spatial risk map, which is a tool that evaluates potential object-occupied spaces, to chart a path that effectively minimizes disruptions. This study presents experiments in which a robot successfully guided a user through the passage of multiple objects and people.
In another study, [65] introduced an egocentric human trajectory forecasting model that was designed for navigation in crowded environments. By leveraging a wearable camera, the model captures detailed surroundings and the user’s past trajectory to predict their future. The trajectories obtained using ORB-SLAM3 were used to train the egocentric human trajectory forecasting model, which predicted the wearer’s future trajectory based on contextual cues captured by the wearable camera and the wearer’s past trajectory.
@@
An intelligent autonomous scooter was developed in [104] for navigating environments with small safety margins and highly dynamic pedestrian traffic such as sidewalks with numerous obstacles and pedestrians. The authors proposed a hybrid mapping solution that combines far-field and near-field mapping to navigate through dynamic environments. This approach utilizes sensor fusion to adapt dynamically to complex and cluttered environments.
@@ III-C2 Practical challenges and considerations
In addition to the technical and methodological aspects, we recognized the importance of practical challenges and considerations that can significantly affect the usability and adoption of SLAM-based assistive systems. Therefore, we included an evaluation of the practical challenges and operational efficiency, as summarized in Tables XXIV and XXV. The information in these tables has either been directly extracted from the article or can be easily inferred from the article’s text. These tables provide information on user-friendliness, cost-efficiency, weight, comfort for extended use, adjustable fit, fatigue mitigation, and portability of the assistive tools described in the reviewed studies. For instance, while smartphones and lightweight devices such as eyeglasses-mounted sensors [74] and ARCore-supported smartphones with haptic gloves [75] are generally well received because of their high portability and ease of use, heavier devices such as guiding robots [68] and rolling suitcase-shaped device [95] are noted to cause user fatigue over extended periods. The augmented cane [67], although found to improve confidence and workload for novice and expert users, also faced usability challenges owing to its weight. Clear instructions and easy learning curves, as seen in electronic glasses with haptic modules [57], play a significant role in enhancing user satisfaction. However, the cost efficiency of these technologies varies, with some solutions being more affordable and accessible.
TABLE XXIV: Practical challenges and operational efficiency - Part I.
@@
TABLE XXV: Practical challenges and operational efficiency - Part II.
@@
This section discusses how the approaches proposed by the studies included in our SLR have the potential to improve navigation for BVI people. These studies focused on diverse attributes, such as accurate pose estimation, semantic mapping, sensor fusion, and algorithmic innovations, to improve the quality of BVI navigation. Table XXVI presents the categorization of attributes that contribute to enhancing the mobility and navigation of visually impaired individuals.
@@ III-D1 Attributes enhancing mobility and navigation
TABLE XXVI: Attributes of SLAM-based navigation systems that contribute to enhancing BVI navigation, along with referenced papers emphasizing each feature.
@@ III-D2 Effectiveness of localization and mapping techniques in real-world scenarios
The effectiveness of the localization and mapping techniques in real-world scenarios varies across studies. Tables XXVII-XXIX summarize these evaluations, highlighting key attributes such as the working area, localization and mapping accuracy level, robustness level, consideration of dynamic objects, and running time. The robustness and accuracy levels reported in these tables are extracted from each paper’s context; each rating reflects conditions specific to that paper and is not necessarily superior or inferior to the other approaches. Thus, these values are not comparable due to differing conditions across the papers.
@@
TABLE XXVII: Effectiveness of localization and mapping techniques in indoor environments as indicated by the literature. (Part I)
@@
TABLE XXVIII: Effectiveness of localization and mapping techniques in indoor environments as indicated by the literature. NM indicates that the information is not explicitly mentioned in the paper. (Part II)
@@
TABLE XXIX: Effectiveness of localization and mapping techniques in outdoor and mixed environments as indicated by the literature. NM indicates that the information is not explicitly mentioned in the paper.
@@ III-D3 User-based evaluations
This section analyzes the user-based evaluations conducted to assess the satisfaction of the proposed SLAM-based assistive systems. By examining these evaluations, we gained insights into the real-world applicability and user acceptance of these technologies. Several studies conducted user-based evaluations with actual participants to assess the effectiveness of and satisfaction with their proposed systems. These evaluations provided valuable insights into the usability and acceptance of assistive technologies. Tables XXX and XXXI summarize studies that include user-based evaluations.
@@
Some studies only conducted technical tests, without involving direct user feedback. These studies are summarized in Table XXXII. For example, [65] and [66] focused on the technical performance of their systems and conducted tests in controlled environments but did not report user satisfaction.
@@
TABLE XXX: User satisfaction evaluation. This table includes studies that conducted user-based evaluations with actual participants to assess the effectiveness and satisfaction of the proposed systems - Part I.
@@
TABLE XXXI: User satisfaction evaluation. This table includes studies that conducted user-based evaluations with actual participants to assess the effectiveness and satisfaction of the proposed systems - Part II.
@@
TABLE XXXII: User satisfaction evaluation. This table includes studies that primarily conducted technical tests without direct user-based evaluations.
@@ III-D4 System prototype information
To provide a comprehensive understanding of the assistive solutions proposed in the reviewed studies, we present the information regarding the system prototypes in Tables XXXIII-XXXVII. These tables include data on the functionalities, sensors used, computing resources, human-computer interaction (HCI) mechanisms, assistive tools, battery life, and whether the solutions are machine learning-based. Notably, the specifications in these tables cover the entire assistive system, and not just the localization and mapping components, as presented in Tables XVII-XIX.
TABLE XXXIII: System prototype information for wearable devices - Part I.
@@
TABLE XXXIV: System prototype information for wearable devices - Part II.
@@
TABLE XXXV: System prototype information for handheld devices - Part I.
@@
TABLE XXXVI: System prototype information for handheld devices - Part II.
@@
TABLE XXXVII: Prototype information for Robot systems, ride-on systems, and setups without specific devices or where device information is not mentioned.
@@
Many systems leverage machine learning for functionalities, such as object detection, scene understanding, and localization. Algorithms such as YOLO, Faster R-CNN, and various deep neural networks are commonly employed. These machine learning-based solutions enhance the accuracy and efficiency of the assistive systems. Table XXXVIII categorizes the machine learning approaches used in assistive devices, along with their references. This categorization illustrates the diversity of machine-learning techniques applied to improve the functionalities of assistive systems for BVI navigation.
TABLE XXXVIII: Categorization of machine learning techniques used in assistive solutions

### v6

@@ Abstract
In recent decades, several assistive technologies have been developed to improve the ability of blind and visually impaired (BVI) individuals to navigate independently and safely. At the same time, simultaneous localization and mapping (SLAM) techniques have become sufficiently robust and efficient to be adopted in developing these assistive technologies. We present the first systematic literature review of 54 recent studies on SLAM-based solutions for blind and visually impaired people, focusing on literature published from 2017 onward. This review explores various localization and mapping techniques employed in this context. We systematically identified and categorized diverse SLAM approaches and analyzed their localization and mapping techniques, sensor types, computing resources, and machine-learning methods. We discuss the advantages and limitations of these techniques for blind and visually impaired navigation. Moreover, we examine the major challenges described across studies, including practical challenges and considerations that affect usability and adoption. Our analysis also evaluates the effectiveness of these SLAM-based solutions in real-world scenarios and user satisfaction, providing insights into their practical impact on BVI mobility. The insights derived from this review identify critical gaps and opportunities for future research activities, particularly in addressing the challenges presented by dynamic and complex environments. We explain how SLAM technology offers the potential to improve the ability of visually impaired individuals to navigate effectively. Finally, we present future opportunities and challenges in this domain.
@@ II-B3 Selection criteria
Table V presents the selection criteria used to identify the eligible studies during the selection process. The Availability criterion included studies accessible in full text from digital databases. In addition, the Language criterion ensured the inclusion of publications written only in English. Furthermore, the Publication Period criterion restricted the inclusion of studies published between January 2017 and July 2023. This timeframe was chosen to prioritize the most current and state-of-the-art approaches in this rapidly evolving field. By focusing on this recent period, we aimed to provide a comprehensive yet manageable review of the latest innovations without overwhelming readers with potentially outdated information. The Type of Source criterion included conference and journal papers, which were considered peer-reviewed and academically recognized sources. Books, dissertations, newsletters, speeches, technical reports, and white papers were excluded. Finally, the Relevance criterion played an important role in the exclusion process; therefore, publications outside the scope of our study were excluded based on a review of their titles and abstracts.
@@ III Result
To further understand the technical features employed in these solutions, the detailed information is presented in Tables XVIIXIX. These tables focus on the localization and mapping components of the assistive system, specifically highlighting the sensor types, computing resources, and application of machine learning-based methods. By examining these features, we can gain deeper insight into how these systems are structured and the diverse technologies utilized to achieve accurate and efficient SLAM for assistive navigation. It is important to note that this information relates only to the localization and mapping components of the assistive navigation solutions. Details of the entire system are provided in Tables XXXIVXXXVIII.
@@ III-A1 Approaches
Dynamic-SLAM Feature-based Optimization-based Visual [90]
@@ III-A2 Sensor type
Table XX provides a detailed breakdown of the sensor types employed in the reviewed studies, highlighting the prevalence of different sensor modalities and their combinations in SLAM-based assistive technologies for BVI individuals.
TABLE XX: Classification of sensor type
Sensor Type References
Camera only [56, 60, 65, 73, 75, 85, 93, 101, 102, 105, 109]
Camera, IMU [98]
RGB-D camera [57, 62, 64, 66, 69, 71, 72, 89, 90, 97]
RGB-D, LiDAR [100]
RGB-D, IMU [70, 74, 80, 84, 88, 91, 92, 96, 99]
LiDAR [61, 67, 68, 112]
LiDAR, IMU [87, 95]
Stereo camera, other sensors [63, 104, 108]
3D time-of-flight [77]
3D time-of-flight, IMU [78, 106]
Fisheye, depth camera [76]
Others [58, 59, 79, 82, 83, 103, 107, 110, 111]
@@ III-A3 Computing resource
To process data and run localization and mapping algorithms, the reviewed studies adopted two classes of computational resources: local and remote. Local computations are performed in situ on devices, such as smartphones, tablets, laptops, portable microcontrollers, and UP board computers. In some cases, algorithms were applied on PCs. Table XXI categorizes the computing resources used in the reviewed studies for localization and mapping tasks. Information regarding the computing resources for the entire navigation assistive system is detailed in Tables XXXVI-XXXVIII
TABLE XXI: Categorization of computing resources used for localization and mapping tasks.
@@
Although SLAM techniques offer significant benefits for navigation across various applications, their use in improving mobility for BVI individuals presents unique considerations. In general applications, the advantages of SLAM include accurate real-time mapping and localization, adaptability to unknown environments, and the ability to function without an external infrastructure. These advantages are essential when applied to BVI navigation. For instance, the ability to provide real-time accurate spatial information is crucial for safe navigation and obstacle avoidance.
Conversely, some limitations of SLAM in general applications, such as computational complexity and sensor dependencies, pose significant challenges in the BVI context. The need for lightweight, portable devices with a long battery life is critical for BVI users, who rely on these systems for extended periods. In addition, robust performance in diverse environments, including crowded spaces and varying lighting conditions, is vital for effective BVI navigation. Unique to BVI applications is the need to translate complex spatial data into intuitive, non-visual feedback. Furthermore, the integration of semantic information to provide context about the environment (e.g., identifying doors, stairs, or pedestrian crossings) in order to interact with the environments during navigation is particularly valuable for BVI users, but may be less critical in other SLAM applications. In the following subsections, we explore in detail the advantages and limitations of SLAM techniques when applied to BVI navigation.
@@ III-B1 Advantages
The cost-effectiveness of SLAM-based solutions attributed to their reliance on widely available low-cost sensors makes this technology particularly interesting. Systems employing monocular cameras or wearable RGB-D cameras exemplify how SLAM can be implemented in a cost-effective manner without compromising functionality, thus making advanced navigational aids accessible to more users [69, 85]. The advantages of SLAM, derived from the literature, are listed in Table XXII.
TABLE XXII: Key features of SLAM that benefit visually impaired navigation.
@@ III-B2 Limitations
Some SLAM applications require external calibration or setup, such as placement of calibration boards in specific environments. This requirement can limit the spontaneity and ease of use of SLAM-based navigation aids because it imposes additional constraints [105]. Table XXIII outlines the overall limitations of SLAM, derived from the publications under review.
TABLE XXIII: Limitations of SLAM for visually impaired navigation.
@@ III-C1 Technical and methodological challenges
Optimal pathfinding, perception of surroundings, and obstacle avoidance are crucial for navigation. Precise localization of a visually impaired user within the environment is essential for the effective operation of these functions. Our surveyed papers addressed the localization and mapping problems using various techniques. Some of these studies explored other navigation-related challenges. We categorized these challenges into two groups: those relevant to environmental complexities, and those related to the sensors used to receive environmental data. Dynamic obstacles and crowded spaces constitute the challenges in the first group, whereas challenges related to changing lighting conditions and the rapid motion of users that results in motion blur fall into the second group. Table XXIV lists studies that investigated these challenges through the integration of SLAM with other approaches. In the following section, we discuss the studies that address these challenges.
TABLE XXIV: Challenges addressed in reviewed studies using SLAM techniques
@@
[59] presented a guide mobile robot engineered for the complexities of navigating different environments while considering dynamic objects and human presence. The robot could handle crowded environments with multiple dynamic objects. To accomplish this, the robot leveraged a spatial risk map, which is a tool that evaluates potential object-occupied spaces, to chart a path that effectively minimizes disruptions. This study presents experiments in which a robot successfully guided a user through the passage of multiple objects and people. The research used Cartography SLAM for off-line mapping. It’s important to note that this paper did not address the dynamic environment through SLAM, but rather used SLAM solely as a tool to pre-build the map of the environment.
In another study, [65] introduced an egocentric human trajectory forecasting model that was designed for navigation in crowded environments. The model predicts the path of the sensor wearer using their past trajectories, nearby pedestrian trajectories, scene semantic and depth data. The authors collected an egocentric human trajectory forecasting dataset. As they could not use GPS or motion capture systems for recording the trajectory, they used ORB-SLAM3 to obtain the ground-truth sensor wearer trajectory. The trajectories obtained using ORB-SLAM3 were used to train the egocentric human trajectory forecasting model. It is important to note that this study, like the previous one, did not handle the crowded environment through SLAM itself, but rather used SLAM as a tool for obtaining the trajectory ground-truth.
@@
An intelligent autonomous scooter was developed in [104] for navigating environments with small safety margins and highly dynamic pedestrian traffic such as sidewalks with numerous obstacles and pedestrians. The authors proposed a hybrid mapping solution that combines far-field and near-field mapping to navigate through dynamic environments. This approach utilizes sensor fusion to adapt dynamically to complex and cluttered environments. However, it should be noted that this study conducted system tests in a completely static environment without moving objects, such as pedestrians. Furthermore, the RTAB-MAP SLAM system was used in this study without any adaptation to dynamic environments.
@@ III-C2 Practical challenges and considerations
In addition to the technical and methodological aspects, we recognized the importance of practical challenges and considerations that can significantly affect the usability and adoption of SLAM-based assistive systems. Therefore, we included an evaluation of the practical challenges and operational efficiency, as summarized in Tables XXV and XXVI. The information in these tables has either been directly extracted from the article or can be easily inferred from the article’s text. These tables provide information on user-friendliness, cost-efficiency, weight, comfort for extended use, adjustable fit, fatigue mitigation, and portability of the assistive tools described in the reviewed studies. For instance, while smartphones and lightweight devices such as eyeglasses-mounted sensors [74] and ARCore-supported smartphones with haptic gloves [75] are generally well received because of their high portability and ease of use, heavier devices such as guiding robots [68] and rolling suitcase-shaped device [95] are noted to cause user fatigue over extended periods. The augmented cane [67], although found to improve confidence and workload for novice and expert users, also faced usability challenges owing to its weight. Clear instructions and easy learning curves, as seen in electronic glasses with haptic modules [57], play a significant role in enhancing user satisfaction. However, the cost efficiency of these technologies varies, with some solutions being more affordable and accessible.
TABLE XXV: Practical challenges and operational efficiency - Part I.
@@
TABLE XXVI: Practical challenges and operational efficiency - Part II.
@@
This section discusses how the approaches proposed by the studies included in our SLR have the potential to improve navigation for BVI people. These studies focused on diverse attributes, such as accurate pose estimation, semantic mapping, sensor fusion, and algorithmic innovations, to improve the quality of BVI navigation. Table XXVII presents the categorization of attributes that contribute to enhancing the mobility and navigation of visually impaired individuals.
@@ III-D1 Attributes enhancing mobility and navigation
TABLE XXVII: Attributes of SLAM-based navigation systems that contribute to enhancing BVI navigation, along with referenced papers emphasizing each feature.
@@ III-D2 Effectiveness of localization and mapping techniques in real-world scenarios
The effectiveness of the localization and mapping techniques in real-world scenarios varies across studies. Tables XXVIII-XXX summarize these evaluations, highlighting key attributes such as the working area, localization and mapping accuracy level, robustness level, consideration of dynamic objects, and running time. The robustness and accuracy levels reported in these tables are extracted from each paper’s context; each rating reflects conditions specific to that paper and is not necessarily superior or inferior to the other approaches. Thus, these values are not comparable due to differing conditions across the papers.
@@
TABLE XXVIII: Effectiveness of localization and mapping techniques in indoor environments as indicated by the literature. (Part I)
@@
TABLE XXIX: Effectiveness of localization and mapping techniques in indoor environments as indicated by the literature. NM indicates that the information is not explicitly mentioned in the paper. (Part II)
@@
TABLE XXX: Effectiveness of localization and mapping techniques in outdoor and mixed environments as indicated by the literature. NM indicates that the information is not explicitly mentioned in the paper.
@@ III-D3 User-based evaluations
This section analyzes the user-based evaluations conducted to assess the satisfaction of the proposed SLAM-based assistive systems. By examining these evaluations, we gained insights into the real-world applicability and user acceptance of these technologies. Several studies conducted user-based evaluations with actual participants to assess the effectiveness of and satisfaction with their proposed systems. These evaluations provided valuable insights into the usability and acceptance of assistive technologies. Tables XXXI and XXXII summarize studies that include user-based evaluations.
@@
Some studies only conducted technical tests, without involving direct user feedback. These studies are summarized in Table XXXIII. For example, [65] and [66] focused on the technical performance of their systems and conducted tests in controlled environments but did not report user satisfaction.
@@
TABLE XXXI: User satisfaction evaluation. This table includes studies that conducted user-based evaluations with actual participants to assess the effectiveness and satisfaction of the proposed systems - Part I.
@@
TABLE XXXII: User satisfaction evaluation. This table includes studies that conducted user-based evaluations with actual participants to assess the effectiveness and satisfaction of the proposed systems - Part II.
@@
TABLE XXXIII: User satisfaction evaluation. This table includes studies that primarily conducted technical tests without direct user-based evaluations.
@@ III-D4 System prototype information
To provide a comprehensive understanding of the assistive solutions proposed in the reviewed studies, we present the information regarding the system prototypes in Tables XXXIV-XXXVIII. These tables include data on the functionalities, sensors used, computing resources, human-computer interaction (HCI) mechanisms, assistive tools, battery life, and whether the solutions are machine learning-based. Notably, the specifications in these tables cover the entire assistive system, and not just the localization and mapping components, as presented in Tables XVII-XIX.
TABLE XXXIV: System prototype information for wearable devices - Part I.
@@
TABLE XXXV: System prototype information for wearable devices - Part II.
@@
TABLE XXXVI: System prototype information for handheld devices - Part I.
@@
TABLE XXXVII: System prototype information for handheld devices - Part II.
@@
TABLE XXXVIII: Prototype information for Robot systems, ride-on systems, and setups without specific devices or where device information is not mentioned.
@@
Table XXXIX categorizes the papers based on the functionalities offered by assistive systems, highlighting the diverse capabilities ranging from basic navigation and obstacle avoidance to advanced features such as scene understanding and social networking. Table XL classifies the papers based on the HCI mechanisms employed, showing the prevalence of audio feedback and the growing trend towards multimodal feedback incorporating haptic and tactile cues. Finally, Table XLI categorizes the studies based on the form factor of the assistive tool, revealing the diversity of approaches, including smartphone-based, wearable devices, handheld devices, and robotic systems.
TABLE XXXIX: Classification of references based on functionalities
Functionality References
Localization [103, 84, 104, 61, 106, 102, 98, 80, 78, 72, 97, 93, 90, 82, 60, 66, 76]
Mapping [104, 107, 64, 90, 66, 69]
Obstacle avoidance [96, 112, 104, 68, 79, 110, 107, 95, 89, 78, 70, 67, 64, 109, 108, 100, 97, 90, 69, 71, 73, 74, 76]
Object detection [102, 89, 83, 87, 77, 72, 64, 93, 82, 63, 73]
Object localization [88, 67, 92, 91, 85, 57, 71]
Scene understanding [100, 97, 82, 58, 63, 66, 71, 74, 75]
Path planning [96, 79, 110, 107, 108, 100, 92, 82, 69, 75]
Face recognition [63, 57]
Remote assistance [58, 63]
Social networking [108, 63]
Way-finding [106, 99, 78, 80, 77, 70, 67, 111, 109, 91, 76, 57]
Dynamic object tracking [90]
Semantic mapping [72, 100]
Collision prediction [95]
Human intent detection [99]
Spatial risk mapping [59]
TABLE XL: Classification of papers based on HCI
HCI Mechanism References
Audio [96, 79, 83, 70, 64, 56, 57, 58, 59, 60, 62, 63, 69, 72, 77, 78, 67, 68, 73, 74, 75, 76, 82, 85, 90, 91, 92, 97, 100, 102, 107, 108, 110, 111, 87, 88, 89, 95, 98, 99, 66]
Tactile [67, 64, 63, 70, 89, 95]
Haptic [57, 75, 68, 107, 88, 91, 92, 111]
Visual hints [76]
Motorized rolling tip [99]
Handle interface [87]
Steering control [104]
Virtual touch [85]
Grounded Kinesthetic [67]
Not Provided/Not mentioned [80, 103, 61, 71, 84, 93, 105, 106, 109, 101]
TABLE XLI: Classification of papers based on assistive tool
Assistive Tool References
Smartphone-based [75, 74, 56, 60, 80, 63, 98, 102, 110, 108]
Glass-based [74, 71, 73, 57, 90, 92, 97, 93, 76]
Cane-based [99, 83, 67, 106, 64, 70, 77, 78, 89, 107, 88]
Robotic systems [59, 68, 79, 87, 95]
Ride-on systems [104, 112]
Wearable sensor [111, 109, 105, 100, 85]
Head-worn sensor [105, 58, 62, 66, 82, 91]
No specific device/ Unspecified [103, 61, 72, 84, 96, 101]
@@
Many systems leverage machine learning for functionalities, such as object detection, scene understanding, and localization. Algorithms such as YOLO, Faster R-CNN, and various deep neural networks are commonly employed. These machine learning-based solutions enhance the accuracy and efficiency of the assistive systems. Table XLII categorizes the machine learning approaches used in assistive devices, along with their references. This categorization illustrates the diversity of machine-learning techniques applied to improve the functionalities of assistive systems for BVI navigation.
TABLE XLII: Categorization of machine learning techniques used in assistive solutions